OpenStudy (firejay5):

Show work! :D Medal will be rewarded! :D 28. Split into a sum of two rational expression with unlike denominators: 2x + 3 over x^2 + 3x + 2

OpenStudy (anonymous):

partial fraction decomposition?

OpenStudy (anonymous):

\[\frac{2x+3}{x^2+3x+2}=\frac{2x+3}{(x+2)(x+1)}\] is a start

OpenStudy (firejay5):

\[\frac{ 2x + 3 }{ x^2 + 3x + 2 }\]

OpenStudy (anonymous):

it is going to be \[\frac{A}{x+1}+\frac{B}{x+2}\] and you need to find \(A\) and \(B\)

OpenStudy (anonymous):

when you add \[\frac{A}{x+1}+\frac{B}{x+2}\] the numerator will be \[A(x+1)+B(x+2)\] and that must be equal to \[2x+3\]

OpenStudy (anonymous):

you can find \(A\) and \(B\) by choosing suitable values for \(x\) in the equation \[A(x+1)+B(x+2)=2x+3\] first choose \(x=-1\) and get \[A(-1+1)+B(-1+2)=2(-1)+3\] or \[B=1\]

OpenStudy (anonymous):

now let \(x=-2\) to find \(A\)

OpenStudy (firejay5):

plug -2 into 2x + 3

OpenStudy (mertsj):

Do what satellite said. He gave you this equation: A(x+1)+B(x+2)=2x+3 and said to let x = -2 to find A